In an Ergonomics lab activity, specifically that for memory recall, 12 students were given ten minutes to try to memorize a list of 20 nonsense words. Each student was then asked to list as many of the words he or she could remember both one hour and twenty-four hours later. The number of words recalled correctly for each student is shown below:
Student A B C D E F G H I J K L
1-hour later 14 9 18 12 13 17 16 16 19 8 15 7
24-hours later 10 6 14 6 8 10 12 10 14 5 10 5
Determine, at the 95% confidence level, that for all such students, the mean number of words recalled after one hour exceeds that recalled after twenty-four hours by 5 words.
1
Expert's answer
2020-05-27T18:52:32-0400
"\def\arraystretch{1.5}
 \begin{array}{c: }
 Student & 1-hour \ later & 24-hours\ later & Difference \\ \hline
 A & 14 & 10 & 4\\
 B & 9 & 6 & 3 \\
 C & 18 & 14 & 4 \\
 D & 12 & 6 & 6 \\
 E & 13 & 8 & 5 \\
 F & 17 & 10 & 7 \\
 G & 16 & 12 & 4 \\
 H & 16 & 10 & 6 \\
 I & 19 & 14 & 5 \\
 J & 8 & 5 & 3 \\
 K & 15 & 10 & 5 \\
 L & 7 & 5 & 2
\end{array}"

"\bar{x_d}=4.5, s_d=1.446, n=12, df=n-1=11"

The following null and alternative hypotheses need to be tested:

"H_0:\mu_d\leq5"

"H_1:\mu_d>5"

This corresponds to a right-tailed test, for which a t-test for one mean, with unknown population standard deviation will be used.

Based on the information provided, the significance level is "\alpha=0.05," and the critical value for a two-tailed test is "t_c=1.796."

The rejection region for this one-tailed test is "R=\{t:t>1.796\}".  

The t-statistic is computed as follows:


"t={\bar{x_d}-\mu_d\over s_d/\sqrt{n}}={4.5-5\over 1.446/\sqrt{12}}\approx-1.198"

Since it is observed that "t=-1.198<1.796=t_c," it is then concluded that the null hypothesis is not rejected. Therefore, there is not enough evidence to claim that the population mean "\mu" is greater than 5, at the 0.05 significance level.

Using the P-value approach: The p-value is "p=0.8719," and since "p=0.8719>0.05," it is concluded that the null hypothesis is not rejected. Therefore, there is not enough evidence to claim that the population mean "\mu" is greater than 5, at the 0.05 significance level.